OpenStudy (superhelp101):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66. 70 72 111 114

OpenStudy (solomonzelman):

\(\large\color{black}{ a_1=-12 }\) and \(\large\color{black}{ a_{27}=-12 }\), So we can say that \(\large\color{black}{ a_{27}=a_1+d(26) }\) Plug in your given values of a1, and a27. \(\large\color{black}{ 66=-12+d(26) }\) can you solve for d ?

OpenStudy (superhelp101):

yes ok thanks can you help me with another one? i'll post it

OpenStudy (solomonzelman):

Hold on, not yet... you need 42nd term.

OpenStudy (superhelp101):

oh yeah that's right...

OpenStudy (solomonzelman):

First, tell me what ` d ` equals.

OpenStudy (superhelp101):

3

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

So after you know that the difference is 3, and a1 is (-12) you can find a42. You can use a1, \(\large\color{black}{ a_{42}=a_{1}+d(42-1) }\) (knowing that a1=-12) OR, you can use a27, \(\large\color{black}{ a_{42}=a_{27}+d(42-27) }\) (knwoing that a27=66)

OpenStudy (superhelp101):

ok so I can use either one?

OpenStudy (solomonzelman):

Oh yeah :)

OpenStudy (superhelp101):

thxx! ;)

OpenStudy (solomonzelman):

tell me what \(\large\color{red}{ a_{42} }\) equals and we are good for this question.

OpenStudy (superhelp101):

give meh just a sec..

OpenStudy (superhelp101):

111

OpenStudy (solomonzelman):

Using \(\large\color{red}{ a_{1} }\) \(\large\color{black}{ a_{42}=a_{1}+d(42-1) }\) \(\large\color{black}{ a_{42}=-12+3(41) }\) \(\large\color{black}{ a_{42}=~~\color\red{??} }\) ────────────────────────── Using \(\large\color{green}{ a_{27} }\) \(\large\color{black}{ a_{42}=a_{27}+d(42-27) }\) \(\large\color{black}{ a_{42}=66+d(15) }\) \(\large\color{black}{ a_{42}=~~\color\green{??} }\)

OpenStudy (solomonzelman):

Yes !

OpenStudy (superhelp101):

Awww thanks for writing all that out, such a help!!! :). Give me like 4 min then I'll post my other ? :)

OpenStudy (solomonzelman):

Sure :)